# PBLean: Pseudo-Boolean Proof Certificates for Lean 4

### v1

@@ Abstract
We present PBLean, a method for importing VeriPB pseudo-Boolean (PB) proof certificates into Lean 4. Key to our approach is reflection: a Boolean checker function whose soundness is fully proved in Lean and executed as compiled native code. Our method scales to proofs with tens of thousands of steps that would exhaust memory under explicit proof-term construction. Our checker supports all VeriPB kernel rules, including cutting-plane derivations and proof-by-contradiction subproofs. In contrast to external verified checkers that produce verdicts, our integration yields Lean theorems that can serve as composable lemmas in larger formal developments. To derive theorems about the original combinatorial problems rather than about PB constraints alone, we support verified encodings. This closes the trust gap between solver output and problem semantics since the constraint translation and its correctness proof are both formalized in Lean. We demonstrate the approach on various combinatorial problems.
@@ 1 Introduction
We present PBLean, a VeriPB integration for Lean 4 that parses kernel-format proofs and verifies them via reflection. Instead of building Lean proof terms for each derivation step, we define a Boolean checker function and prove its soundness once; Lean’s native_decide then runs the checker as compiled code. This follows the pattern of Lean’s bv_decide tactic [21] and scales to instances that would exhaust memory under explicit proof construction. Our implementation covers all VeriPB kernel rules, including proof-by-contradiction subproofs for optimization reasoning. We apply the verified encoding workflow to independence numbers of Paley graphs and other combinatorial theorems to obtain end-to-end-verified results.
@@ 2 Background
The VeriPB proof format [10, 4] provides two modes: an augmented mode with high-level rules (symmetry breaking, redundance-based strengthening) and a kernel mode (version 3.0) with only primitive operations. The kernel format’s small rule set of just six rule types enables a rigorous soundness proof and makes it well-suited for theorem prover integration. It supports: pol (polynomial operations including all cutting-planes rules), rup (reverse unit propagation with conflict hints), pbc (proof by contradiction with inline subproofs), del (constraint deletion, sound by monotonicity), sol/soli (solution logging), and conclusion (UNSAT/SAT/BOUNDS verdict). Solvers such as RoundingSat produce augmented proofs; veripb --elaborate translates them to kernel format.
@@ 3 Lean Formalization
Our 13 soundness lemmas are summarized in Table 1; every lemma is proved in Lean and verified by the kernel. Two details matter for matching VeriPB’s semantics: normalization must precede saturation and division, and the negation of $\sum a_{i}\ell_{i}\geq d$ is $\sum a_{i}\bar{\ell}_{i}\geq(\sum a_{i})-d+1$.
@@ 4 Implementation
PBLean is a standalone Lean 4 project (v4.28.0) with no Mathlib dependency. The kernel layer (PseudoBoolean.lean, $\scriptstyle\sim$700 lines) consists of types and 13 soundness lemmas. The code is available on GitHub. [^1] https://github.com/leansolving/pblean
The reflection-based checker (Reflect.lean, $\scriptstyle\sim$1800 lines) uses a line-by-line tokenizer to convert VeriPB proofs into an abstract syntax tree, handling nested pbc subproofs recursively. It maintains a constraint database organized as a hash map, with an @[implemented_by] annotation providing an array-based runtime replacement for performance without weakening the proof. Each rule updates or queries this database: pol evaluates the RPN expression, rup derives a conflict by negating the target and applying PB unit propagation, and pbc saves the database, adds the negated assumption, verifies inner steps recursively, then restores the snapshot. The checker returns true only when a contradictory constraint arises, and we establish soundness by induction over proof steps, connecting the Boolean computation to the kernel lemmas.
@@
Figure 1: End-to-end workflow for trusted encodings. OPB (Optimized Pseudo-Boolean) is the standard input format for PB solvers. Solid boxes are verified in Lean; the dashed box is untrusted. The dotted arrow represents the bridge theorem composing encoding correctness with the PB refutation.
@@ 5 Evaluation
The proofs rely on cutting-planes (pol) and reverse unit propagation (rup) steps; the largest instance (2525 edges, about 63 000 proof lines) verifies in about 3.5 minutes.
@@
As an unexpected discovery worth noting, the data yields a formally verified proof that $\alpha(\mathrm{Paley}(p))$ is not monotone in $p$: $\alpha(\mathrm{Paley}(97))=6>5=\alpha(\mathrm{Paley}(101))$. Here, each equality combines the VeriPB upper bound with a lower bound from an explicit witness verified by native_decide.
@@
Beyond independence numbers of Paley graphs, we provide trusted encodings for five additional combinatorial problems. These applications demonstrate the generality of the verified encoding approach. The underlying theorems are well known; the contribution lies in the end-to-end formal verification within Lean. Table 3 reports the same pipeline metrics as Table 2.
Table 3: Additional verified combinatorial results, using the reflection-based checker. Column headers as in Table 2.
@@
The impossibility of $L(2,6)$ is proved via the reflection checker from a 223-line kernel proof (72 variables, 463 constraints).
@@
The lower bound uses the witness coloring $\{1,4\}\mapsto\mathrm{true}$, $\{2,3\}\mapsto\mathrm{false}$. The upper bound ($\{1,\ldots,5\}$ has no Schur-free coloring) is verified via the reflection checker from a 13-line kernel proof.
@@
The lower bound uses the witness RRBBRRBB; the upper bound ($\{1,\ldots,9\}$ has no AP-free coloring) is verified via the reflection checker in a 19-line kernel proof.
@@
The lower bound uses the pentagon ($C_{5}$) coloring of $K_{5}$. The upper bound—that $K_{6}$ admits no Ramsey-free coloring—is verified via the reflection checker from a 31-line kernel proof derived from 15 edge variables and 40 constraints (20 triangles).
@@ 6 Related Work
CakePB [13] provides end-to-end verified subgraph solving in HOL4/CakeML, including both a PB proof checker and verified graph-problem encoders compiled to a standalone binary. As discussed in Section 1, the approaches are complementary: CakePB targets standalone verification with a small trusted computing base, while PBLean yields Lean theorems that can serve as lemmas in larger formal developments. Several extensions of the VeriPB framework are known, certifying MaxSAT preprocessing [18], solution-improving search [1], dynamic programming and decision diagrams [8], and pseudo-Boolean optimization [19].

### v2

@@ Abstract
We present PBLean, a method for importing VeriPB pseudo-Boolean (PB) proof certificates into Lean 4. Key to our approach is reflection: a Boolean checker function whose soundness is fully proved in Lean and executed as compiled native code. Our method scales to proofs with tens of thousands of steps that would exhaust memory under explicit proof-term construction. Our checker supports all VeriPB kernel rules, including cutting-plane derivations, proof-by-contradiction subproofs, and redundance-based reasoning for symmetry breaking. In contrast to external verified checkers that produce verdicts, our integration yields Lean theorems that can serve as composable lemmas in larger formal developments. To derive theorems about the original combinatorial problems rather than about PB constraints alone, we support verified encodings. This closes the trust gap between solver output and problem semantics since the constraint translation and its correctness proof are both formalized in Lean. We demonstrate the approach on various combinatorial problems.
@@ 1 Introduction
We present PBLean, a VeriPB integration for Lean 4 that parses kernel-format proofs and verifies them via reflection. Instead of building Lean proof terms for each derivation step, we define a Boolean checker function and prove its soundness once; Lean’s native_decide then runs the checker as compiled code. This follows the pattern of Lean’s bv_decide tactic [21] and scales to instances that would exhaust memory under explicit proof construction. Our implementation covers all VeriPB kernel rules, including proof-by-contradiction subproofs for optimization reasoning and redundance/dominance rules for symmetry breaking. We apply the verified encoding workflow to independence numbers of Paley graphs and other combinatorial theorems to obtain end-to-end-verified results.
@@ 2 Background
The VeriPB proof format [10, 4] provides two modes: an augmented mode with high-level rules (symmetry breaking, redundance-based strengthening) and a kernel mode (version 3.0) with only primitive operations. The kernel format’s small rule set of just eight rule types enables a rigorous soundness proof and makes it well-suited for theorem prover integration. It supports: pol (polynomial operations including all cutting-planes rules), rup (reverse unit propagation with conflict hints), pbc (proof by contradiction with explicit subproofs), red/dom (redundance and dominance, adding equisatisfiable constraints via a substitution witness), del (constraint deletion, sound by monotonicity), weaken (coefficient weakening), sol/soli (solution logging), and conclusion (UNSAT/SAT/BOUNDS verdict). Solvers such as RoundingSat produce augmented proofs; veripb --elaborate translates them to kernel format.
@@ 3 Lean Formalization
Our 15 soundness lemmas are summarized in Table 1; every lemma is proved in Lean and verified by the kernel. Two details matter for matching VeriPB’s semantics: normalization must precede saturation and division, and the negation of $\sum a_{i}\ell_{i}\geq d$ is $\sum a_{i}\bar{\ell}_{i}\geq(\sum a_{i})-d+1$.
@@
Redundance (red/dom)
applySubstConstr_sat_rev $C|_{\omega}.\mathit{sat}\ v\to C.\mathit{sat}\ (\omega(v))$
constr_sat_noSubst $\mathit{unaffected}(\omega,C)\to C.\mathit{sat}\ v\to C.\mathit{sat}\ (\omega(v))$
@@
Trust model.
The reflection-based checker relies on Lean.trustCompiler, the axiom that Lean’s compiled native code agrees with its kernel evaluator. This is the same trust assumption underlying bv_decide, native_decide, and omega. Lean’s kernel does not verify native compilation; a bug in the compiler could in principle produce an incorrect evaluation result. In practice, the Lean compiler is mature and widely used, and the alternative—explicit proof-term construction—does not scale beyond small instances. The explicit proof-term approach (our first method above) avoids this axiom entirely, providing a fallback for instances where Lean.trustCompiler is unacceptable.
@@ 4 Implementation
PBLean is a standalone Lean 4 project (v4.28.0-rc1) with no Mathlib dependency. The kernel layer (PseudoBoolean.lean, $\scriptstyle\sim$900 lines) consists of types and 15 soundness lemmas. The code is available on GitHub. [^1] https://github.com/leansolving/pblean
The reflection-based checker (Reflect.lean, $\scriptstyle\sim$2500 lines) uses a line-by-line tokenizer to convert VeriPB proofs into an abstract syntax tree, handling nested pbc subproofs recursively. It maintains a constraint database organized as a hash map, with an @[implemented_by] annotation providing an array-based runtime replacement for performance without weakening the proof. Each rule updates or queries this database: pol evaluates the RPN expression, rup derives a conflict by negating the target and applying PB unit propagation, pbc saves the database, adds the negated assumption, verifies inner steps recursively, then restores the snapshot, and red/dom verify coverage of affected constraints and process subproofs for each proof goal before adding the equisatisfiable constraint. The checker returns true only when a contradictory constraint arises, and we establish soundness by induction over proof steps, connecting the Boolean computation to the kernel lemmas.
@@
Figure 1: End-to-end workflow for trusted encodings. OPB is the standard input format for PB solvers. Solid boxes are verified in Lean; the dashed box is untrusted. The dotted arrow represents the bridge theorem composing encoding correctness with the PB refutation.
@@ 5 Evaluation
The proofs rely on cutting-planes (pol) and reverse unit propagation (rup) steps; the largest instance (2526 edges, about 63 000 proof lines) verifies in about 3.5 minutes. The pigeonhole example in Table 3 additionally exercises the red rule for symmetry breaking.
@@
A noteworthy byproduct of this evaluation is a formally verified proof that $\alpha(\mathrm{Paley}(p))$ is not monotone in $p$: $\alpha(\mathrm{Paley}(97))=6>5=\alpha(\mathrm{Paley}(101))$. Here, each equality combines the VeriPB upper bound with a lower bound from an explicit witness verified by native_decide.
@@
Beyond independence numbers of Paley graphs, we provide trusted encodings for seven additional combinatorial problems. These applications demonstrate the generality of the verified encoding approach. The underlying theorems are well known; the contribution lies in the end-to-end formal verification within Lean. Table 3 reports the same pipeline metrics as Table 2.
Table 3: Additional verified combinatorial results, using the reflection-based checker. Column headers as in Table 2. Dashes indicate a hand-crafted proof (no solver call).
@@
[…]
The impossibility of $L(2,6)$ is proved from a 223-line kernel proof. The larger instance $L(2,9)$ (162 variables, 1 867 constraints) requires a 44 342-line proof and takes about 155 s to verify, demonstrating that the approach scales to proofs with tens of thousands of steps.
@@
[…]
The $S(2)$ lower bound uses the witness coloring $\{1,4\}\mapsto\mathrm{true}$, $\{2,3\}\mapsto\mathrm{false}$. The upper bound ($\{1,\ldots,5\}$ has no Schur-free coloring) is verified from a 17-line kernel proof. For $S(3)=13$, the encoding generalizes to $k$-colorings using a one-hot representation: $k$ binary variables per element plus at-least-one and no-monochromatic-triple constraints per color. The resulting instance (42 variables, 161 constraints) is verified from a 209-line proof.
@@
[…]
The $W(2,3)$ lower bound uses the witness RRBBRRBB; the upper bound ($\{1,\ldots,9\}$ has no AP-free coloring) is verified from a 19-line kernel proof. The generalized encoding supports $k$-term arithmetic progressions. For $W(2,4)=35$, the upper bound ($\{1,\ldots,35\}$ has no 2-coloring avoiding monochromatic 4-APs) is verified from a 429-line proof over 36 variables.
@@
theorem ramsey9_34_impossible :
[…]
The $R(3,3)$ lower bound uses the pentagon ($C_{5}$) coloring of $K_{5}$. The upper bound—that $K_{6}$ admits no Ramsey-free coloring—is verified from a 31-line kernel proof. The generalized encoding supports asymmetric Ramsey numbers $R(s,t)$: forbid red $K_{s}$ and blue $K_{t}$ separately. For $R(3,4)\leq 9$, the instance has 36 edge variables and 210 constraints (84 triangle and 126 $K_{4}$ constraints); the 25 194-line proof verifies in about 26 s.
@@
Pigeonhole principle.
The pigeonhole principle $\mathrm{PHP}(m,n)$ states that $m$ pigeons cannot be placed into $n$ holes with at most one pigeon per hole when $m>n$. The proof of $\neg\mathrm{PHP}(3,2)$ uses the red rule to add a symmetry-breaking constraint (without loss of generality, pigeon 1 goes to hole 1) via a cyclic substitution witness. This is the only example exercising the redundance rule.
theorem php32_red_reflect : formulaUnsat
Bin packing.
Given $n$ items with sizes $s_{0},\ldots,s_{n-1}$ and $m$ bins of capacity $C$, the bin packing problem asks whether all items can be assigned to bins without exceeding capacity. The capacity constraint $\sum_{i}s_{i}\cdot x_{i,j}\leq C$ is a single PB inequality; in CNF it requires a sequential weight counter encoding [24] with $O(n\cdot C)$ auxiliary variables per bin. Our instance has 12 items (sizes $[10,9,8,8,6,5,4,4,4,4,4,4]$), 5 bins, and $C=14$. The total size equals $5\times 14=70$, so total weight alone does not refute feasibility. The instance is not AMO-reducible: pairs like $10+4=14$ and triples like $4+4+4=12$ fit in a single bin. Unsatisfiability follows from a rounding argument: dividing each capacity constraint by 4 shows that each bin holds at most 3 “slots,” heavy items ($\geq 8$) consume 2 slots and light items ($\geq 4$) consume 1 slot, giving a total demand of $2\times 4+8=16>15=3\times 5$. The PB encoding consists of just 17 constraints over 60 variables; the VeriPB kernel proof has 687 lines. A CNF encoding of the same instance via the standard sequential weight counter [24] requires 960 variables (900 auxiliary) and 2 617 clauses, and this encoding overhead grows linearly with the capacity $C$.
[…]
@@ 6 Related Work
CakePB [13] provides end-to-end verified subgraph solving in HOL4/CakeML, including both a PB proof checker and verified graph-problem encoders compiled to a standalone binary. As discussed in Section 1, the approaches are complementary: CakePB targets standalone verification with a small trusted computing base, while PBLean yields Lean theorems that can serve as lemmas in larger formal developments. Running CakePB on our Paley instances, it checks $\mathrm{Paley}(101)$ in under 1 s; PBLean takes about 200 s. This roughly $200\times$ overhead is the cost of producing a Lean theorem rather than a standalone verdict. Several extensions of the VeriPB framework are known, certifying MaxSAT preprocessing [18], solution-improving search [1], dynamic programming and decision diagrams [8], and pseudo-Boolean optimization [19].